There is a uniformly charged non conducting solid sphere made up of material of dielectric constant one. If electric potential at infinity be zero, then the potential at its surface is V. If we take electric potential at its surface to be zero then the potential at the center be 1. 3V/22.V3.V/24.Zero

Grade:12

3 Answers

Vikas TU
6 years ago
Dear Student,
There is a uniformly charged non conducting solid sphere made up of material of dielectric constant one. If electric potential at infinity be zero, then the potential at its surface is V.
For non conducting sphere, the potential at its surface is equal to potential at center.
And here is given that potential at its surface is zero then at center potential is also zero.
 
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)
Venkatesh
4 years ago
for uniformly charged volumes the pot at centre is 3/2 times the surface
Pot at centre vc =3/2v
Pot at surface=vsa
Vc-vs=3/2v-v=v/2
Rishi Sharma
3 years ago
Dear Student,
Please find below the solution to your problem.

for uniformly charged volumes the pot at centre is 3/2 times the surface
Pot at centre vc =3/2v
Pot at surface=vsa
Vc-vs=3/2v-v=v/2

Thanks and Regards